OpenStudy (anonymous):

Eliminate the "z" variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1

OpenStudy (anonymous):

Multiply 2x-2y-z by 2. You'll get: 4x-4y-2z Now adding the last two equations; \[(2x-y+2z+15)+(4x-4y-2z-2)= 6x-5y+13\]

OpenStudy (anonymous):

You can also write it as: 6x-5y= -13

OpenStudy (anonymous):

*Sorry. Correction. You'll get 4x-4y-2z=2